A rectangle is 2 times as long as it is wide, and its perimeter is 54 centimeters. Find its dimensions.
Width _______cm
Length_______cm

Answer by addingup(3677)

L = 2W
2L + 2W = 54
2(2W) + 2W = 54
4W + 2W = 54
6W = 54
W = 9 this is the width. Since L = 2W:
L = 2(9) = 18 this is your length